OpenStudy (anonymous):

what is the probability that robert will draw a blue marble and then a green marble if he does not replace the first marble?

OpenStudy (anonymous):

options are- 2/15 7/10 3/25 22/45

OpenStudy (anonymous):

how many marbles is the blue one and green one

OpenStudy (anonymous):

it doesnt say

OpenStudy (anonymous):

well yea its red 2 blue 4 green 3 and orange 1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

please help me out im on a bit of a time crunch

OpenStudy (anonymous):

okay im thinking

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

7/10

OpenStudy (anonymous):

thanks can i ask you a few more questions

OpenStudy (anonymous):

its 7/10 because if you add up all the colcors their is tena nd you add the ble and green marble beacuse what eer you do to one side you do to the other

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

robert has these marbles in a jar. of robert draws one marble without looking. what is the probability that it will be green? it nuses that same graph as the one above

OpenStudy (anonymous):

_uses

OpenStudy (anonymous):

2 + 4 + 3 + 1 = your denominator 3 + 4 = blue and green/ your numerator

OpenStudy (anonymous):

thanks tousii45

OpenStudy (anonymous):

Yup(:

OpenStudy (anonymous):

its for one green right

OpenStudy (anonymous):

the probability it was green is 3/10

OpenStudy (anonymous):

3/10 is your answer if you count all 3

OpenStudy (anonymous):

To draw blue\[ \frac 4{10} \]To drag green after \[ \frac 39 \]Both probability \[ \frac 4{10}\times \frac 39 = \frac{12}{90} = \frac{6}{45} = \frac{2}{15} \]

OpenStudy (anonymous):

maria has a spinner numbered from 1 to 8. if she spins it twice in a row, what is the probability thazt she will spin a 2 and then a 5?

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

2/8

OpenStudy (anonymous):

its one of them im just not sure

OpenStudy (anonymous):

options are 1/64 1/56 14 11/15

OpenStudy (anonymous):

hold on.....

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

i would say 1/64 i learning this now

OpenStudy (anonymous):

multiply 1/8 times 1/8 and you get 1 ad 64

OpenStudy (anonymous):

thanks your the best i just have 2 more questions if you dont mind?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

in problem 4 , wich best describes the two spins of the spinner? dependent events independent events predictable events replacable events

OpenStudy (anonymous):

r u good at ratios ??

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

i would say it is dependent or replaceable. yeah its dependent

OpenStudy (anonymous):

thanks everybody youv been very helpful

OpenStudy (anonymous):

welcome